# New simple and accurate approximations for quintessence

### v1

@@ Abstract
We derive new simple approximations for quintessence in a spatially homogeneous, isotropic and flat spacetime. The approximations are typically an order of magnitude more accurate than anything available in the literature, which from an observational perspective makes numerical calculations superfluous. For example, our tracking quintessence approximation, with one less parameter than CPL and a comparably simple expression, yields a $0.1\%$ maximum relative error of $H(z)/H_{0}$ for the inverse power law scalar field potential. We situate the new results in a broader dark energy context, where the present parametrized deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
@@ 1 Introduction
Some DE models, however, endeavour to have some theoretical motivation, where one of the simplest attempts is quintessence, i.e. a canonical scalar field, $\phi$, that is minimally coupled to gravity, see e.g. [28] and [29] for references. This is the main topic of the present paper, which deals with a non-interacting scalar field and matter in a spatially homogeneous, isotropic and flat spacetime, although we note that the ideas in this paper are applicable to many other models and alternative theories. More specifically, we will derive simple and accurate approximations for thawing and tracking quintessence, [^1] According to Tsujikawa (2013) [28] there are three types of quintessence in the literature: thawing, tracking and scaling freezing quintessence, see [30] for subtleties. In this paper we consider tracking quintessence for asymptotically inverse power law scalar field potentials $V\propto\phi^{-p}$, $p>0$, when $\phi\rightarrow 0$. Since scaling freezing quintessence does not yield past continuous deformations of $\Lambda$CDM we will only briefly touch on this topic in appendix B. which yield continuous deformations of $\Lambda$CDM cosmology.
The outline of the paper is as follows. Motivated by our new quintessence approximations, the next section provides a general DE context. In Section 3 we derive the expressions for the parameters appearing in the previous DE section for thawing and tracking quintessence. The new thawing and tracking quintessence approximations are compared with numerical calculations and previous results in the literature in section 4. The main part of the paper ends with section 5, which includes suggestions for future applications of the present approach and results. Appendix A reviews quintessence approximations found in the literature, while appendix B deals with scaling freezing quintessence.
@@ 2.1 Preliminaries
To find interesting alternatives to $\Lambda$CDM some authors simply choose some $H(z)$, e.g. $H^{2}$ being a polynomial in $1+z$ as in [27], or obtain $H(z)$ by assuming a specific form for one of the distance measures, e.g. $D_{L}(z)$, which yield $H(z)$ via $d[D_{L}(z)/(1+z)]/dz=1/H(z)$, see e.g. [27]. More commonly, however, is to make a choice of the DE equation of state parameter $w_{\mathrm{DE}}(z)=p_{\mathrm{DE}}(z)/\rho_{\mathrm{DE}}(z)$ that is different than the $\Lambda$CDM value $w_{\mathrm{DE}}=-1$ (see e.g. [23] and references therein for a plethora of examples of $w_{\mathrm{DE}}(z)$ and $w_{\mathrm{DE}}(a)$). Note, however, that $w_{\mathrm{DE}}$ is not a physical observable. Since cosmological observations depend on $H(z)$, it is necessary to compute $H(z)$ from $w_{\mathrm{DE}}(z)$, which requires the Einstein field equations and the matter conservation equations.
@@ 2.2 Relationships between $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$, $w_{\mathrm{DE}}$ and $H$
Next we introduce the dimensionless and bounded Hubble normalized variables $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$ and the dimensionless DE equation of state parameter $w_{\mathrm{DE}}$:
$\Omega_{\mathrm{DE}}=\frac{\rho_{\mathrm{DE}}}{3H^{2}}=\frac{\rho_{\mathrm{DE}}}{\rho},\qquad\Omega_{\mathrm{m}}=\frac{\rho_{\mathrm{m}}}{3H^{2}}=\frac{\rho_{\mathrm{m}}}{\rho},\qquad w_{\mathrm{DE}}=\frac{p_{\mathrm{DE}}}{\rho_{\mathrm{DE}}},$ (7)
@@
If one instead of $w_{\mathrm{DE}}(N)$ has $\Omega_{\mathrm{DE}}(N)$, then one can obtain $w_{\mathrm{DE}}(N)$ and $E_{\mathrm{DE}}^{2}(N)$, since (5) and (7) lead to
@@
$E_{\mathrm{DE}}^{2}$ $=\exp(-3N)\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}}}\right)=\exp(-3N)\left(\frac{1-\Omega_{\mathrm{DE}0}}{1-\Omega_{\mathrm{DE}}}\right).$ (13b)
@@ 2.4 Past DE series expansions
which is indeed the case for the quintessence models. To improve the approximation we now perform series expansions in
@@
where $T\rightarrow 0$ when $N\rightarrow-\infty$.
For brevity we only consider $w_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}$. If one starts with a series expansion in $w_{\mathrm{DE}}$ and uses (6) and (12) one obtain an expression for $\Omega_{\mathrm{DE}}$ that subsequently can be series expanded in $T$; on the other hand, if one starts with a series expansion in $\Omega_{\mathrm{DE}}$ and inserts this into (13a) one obtains an expression for $w_{\mathrm{DE}}$ that one then can series expand in $T$. These two procedures result in the same series expansions, which can be written as [^5] The given expressions can be generalized by writing $w_{\mathrm{DE}}=w_{\infty}\left[1+\alpha T+\sigma T^{2}+\dots\right]$, which subsequently can be used to compute the series expansions for $\Omega_{\mathrm{DE}}$ and, e.g., $E_{\mathrm{DE}}^{2}$. The reason we have written $\alpha$ as $1-\gamma$ and $\sigma$ as $(\gamma-1)\beta$ is that this simplifies and is compatible with our thawing and tracking quintessence approximations, where these use the $[1/1]$ Padé expression for $\Omega_{\mathrm{DE}}$ in order to connect our approximations to the present time.
@@
$\Omega_{\mathrm{DE}}$ $=T\left(1-\gamma T+\left[1+\frac{1}{2}(\gamma-1)\left(3+\gamma+\beta\right)\right]T^{2}+\dots\right),$ (21b)
@@ 2.5 DE Padé approximants
To connect with the present time we can set $N=0$ and solve for $T=T_{0}$ in (22) in terms of the left hand side $w_{\mathrm{DE}0}$. However, since $w_{\mathrm{DE}0}$ is not an observable we will instead relate $T_{0}$ to $\Omega_{\mathrm{DE}0}$ and thereby obtain a direct connection to the observables $H_{0}$ and $\rho_{\mathrm{m}0}$ since $1-\Omega_{\mathrm{DE}0}=\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$. To obtain accurate approximations at the present time for $\Omega_{\mathrm{DE}}$ we consider the Padé approximant
@@
This is then inserted into (22a), which when used in (6) to obtain $f(N)$ leads to the same expression as (23) with (24) inserted. To obtain a better approximation for $\Omega_{\mathrm{DE}}(N)$ one can then go on to Padé approximants also involving $\beta$ in the expansion. [^6] These are given by $[1/2]_{\Omega_{\mathrm{DE}}}=\frac{T}{1+\gamma T-\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}$ and $[2/1]_{\Omega_{\mathrm{DE}}}=\frac{\gamma T+\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}{\gamma+\left[1+\frac{1}{2}(\gamma-1)\left(3+\beta+\gamma\right)\right]T}$. To express $T_{0}$ in terms of $\Omega_{\mathrm{DE}0}$ obtained in this manner requires solving quadratic equations in $T_{0}$, which leads to rather messy expressions. The approximation (23) turns out to be a fairly good approximations for quintessence. For this reason we will use the simple result in (24) also for $T_{0}$ in (22b). This leads to expressions for $w_{\mathrm{DE}}$ that can be written as follows: [^7] Note that one can use the series expansion in $T$ for $\Omega_{\mathrm{DE}}$ to obtain the following expression for $E_{\mathrm{DE}}^{2}$: $E_{\mathrm{DE}}^{2}\approx\Omega_{\mathrm{m}0}\exp(-3N)[1+T-(\gamma-1)T^{2}(1+{\textstyle{1\over 2}}(\gamma-1+\beta)T)+\dots]$ and use (24) for $T_{0}$. This expression can subsequently be used to obtain various Padé approximants to obtain good quintessence approximations for $E_{\mathrm{DE}}=E_{\phi}$ and then use $E_{\phi}(N)$ to compute, e.g., $w_{\mathrm{DE}}=w_{\phi}$, if one is so inclined.
@@
which lead to the following present time values
@@
Note that setting $\beta=\gamma-1$ in $[1/1]_{w_{\mathrm{DE}}}(T)$ yields $[0/1]_{w_{\mathrm{DE}}}(T)$ and that setting $\gamma=1$ in both $[1/1]_{w_{\mathrm{DE}}}(T)$ and $[0/1]_{w_{\mathrm{DE}}}(T)$ results in $w_{\mathrm{DE}}=w_{\infty}$, i.e., $\gamma=1$ leads to the $w$CDM model when $w_{\infty}\in(-1,0)$ and the $\Lambda$CDM model when $w_{\infty}=-1$.
@@ 2.6 The Chevallier-Polarski-Linder parametrization
The so-called CPL parameters [25, 26] are often used in cosmology and parameterize $w_{\mathrm{DE}}$ according to
@@
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{(1-\Omega_{\mathrm{DE}0})^{2}},$ (32a)
@@
One can regard (30) as a series expansion in $e^{N}$, but this is inconsistent with the past matter dominant requirement $\Omega_{\mathrm{DE}}\rightarrow 0$ and that one of the CPL parameters can be interpreted as $w_{\infty}$, since this naturally leads to an expansion in $e^{-3w_{\infty}N}$, which is particularly pertinent as observations probe increasingly large redshift. CPL thus mixes this asymptotically motivated function (arising from the $w_{\infty}$ part in (30)) in $\Omega_{\mathrm{CPL}}$ and $E_{\mathrm{CPL}}^{2}$ with another function that comes from the $w_{a}$ part in (30). Due to this mixture of functions, especially the exponential part involving $\frac{z}{1+z}$, the CPL parametrization arguably does not lead to simpler observables than our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based results, and it certainly is not a good approximation for thawing and tracking quintessence.
@@ 3.1 Thawing quintessence
To obtain a system that is useful for deriving expressions for $\gamma$ and $\beta$ for thawing quintessence [^9] The concepts of thawing and freezing were defined by Caldwell and Linder (2005) [32] as follows: thawing is characterized by $w_{\varphi}\approx-1$ where $w_{\varphi}$ subsequently grows, i.e. $w_{\varphi}^{\prime}>0$, while $w_{\varphi}>-1$ and $w_{\varphi}^{\prime}<0$ holds for freezing. we introduce the following dimensionless bounded variable [^10] The variable $\Sigma_{\phi}$ was first introduced by Coley et al. (1997) [33] and Copeland et al. (1998) [34] whose $x$ is $\Sigma_{\varphi}$. Since then, $\Sigma_{\phi}$ (or $\phi^{\prime}$) is commonly used to describe scalar fields in cosmology, see, e.g., Urena-Lopez (2012) [35], equation (2.3), Tsujikawa (2013) [28], equation (16) and Alho and Uggla (2015) [36], equation (8). The reason for using the notation $\Sigma$ for the kernel is because $\Sigma_{\phi}$ plays a similar role as Hubble-normalized shear, which is typically denoted with the kernel $\Sigma$, see e.g. [37].
@@
as follows from the definitions. In addition, the (non-linear) Klein-Gordon equation $\ddot{\varphi}=-3H\dot{\varphi}-V_{,\varphi}$, expressed in $e$-fold time, is needed. This leads to the dynamical system

### v2

@@ Abstract
We derive new simple approximations for quintessence solutions to Einsteins’s field equations in spatially homogeneous, isotropic and flat spacetimes. The approximations are typically an order of magnitude more accurate than anything available in the literature, which from an observational perspective makes numerical calculations superfluous. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We situate the new results in a broader dark energy context, where the present parametrized deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
@@ 1 Introduction
Some DE models, however, endeavour to have some theoretical motivation, where one of the simplest attempts is quintessence, i.e. a canonical scalar field, $\phi$, that is minimally coupled to gravity, see e.g. [28] and [29] for references. This is the main topic of the present paper, which deals with a non-interacting scalar field and matter in a spatially homogeneous, isotropic and flat spacetime, although note that the ideas in this paper are applicable to many other models and alternative theories. More specifically, we will derive simple and accurate approximations for thawing and tracking quintessence solutions, [^1] According to Tsujikawa (2013) [28] there are three types of quintessence in the literature: thawing, tracking and scaling freezing quintessence, see [30] for subtleties. In this paper we consider tracking quintessence for asymptotically inverse power law scalar field potentials $V\propto\phi^{-p}$, $p>0$, when $\phi\rightarrow 0$. Since scaling freezing quintessence does not yield past continuous deformations of $\Lambda$CDM we will only briefly touch on this topic in appendix B. which yield continuous deformations of $\Lambda$CDM cosmology. Notably, the approximations are typically simpler and an order of magnitude more accurate than anything available in the literature. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We note that from an observational perspective the remarkable accuracy of our approximations makes numerical calculations superfluous when assessing observational constraints; moreover the approximations yield an analytic overview of the observational viability of broader ranges of models than those that are typically numerically investigated. For example, it would be interesting to replace the recent numerical observational constraint analysis by Shlivko and Steinhardt [31] with our thawing quintessence approximations for a broader set of models, although we note that our new parametrized DE deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
The outline of the paper is as follows. Motivated by our new quintessence approximations, the next section provides an associated general DE context. In Section 3 we derive the expressions for the parameters appearing in the previous DE section for thawing and tracking quintessence. The new thawing and tracking quintessence approximations are compared with numerical calculations and previous results in the literature in section 4. The main part of the paper ends with section 5, which includes suggestions for future applications of the present approach and results. Appendix A reviews quintessence approximations found in the literature, while appendix B deals with scaling freezing quintessence.
@@ 2.1 Preliminaries
To find interesting alternatives to $\Lambda$CDM some authors simply choose some $H(z)$, e.g. $H^{2}$ being a polynomial in $1+z$ as in [27], or obtain $H(z)$ by assuming a specific form for one of the distance measures, e.g. $D_{L}(z)$, which yield $H(z)$ via $d[D_{L}(z)/(1+z)]/dz=1/H(z)$, see e.g. [27]. More commonly, however, is to make a choice of the DE equation of state parameter
$w_{\mathrm{DE}}=\frac{p_{\mathrm{DE}}}{\rho_{\mathrm{DE}}}$ (4)
that is different than the $\Lambda$CDM value $w_{\mathrm{DE}}=-1$ (see e.g. [23] and references therein for a plethora of examples of $w_{\mathrm{DE}}(z)$ and $w_{\mathrm{DE}}(a)$). Note, however, that $w_{\mathrm{DE}}$ is not a physical observable. Since cosmological observations depend on $H(z)$, it is necessary to compute $H(z)$ from $w_{\mathrm{DE}}(z)$, which requires the Einstein field equations and the matter conservation equations.
@@ 2.2 Relationships between $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$, $w_{\mathrm{DE}}$ and $H$
Next we introduce the dimensionless and bounded Hubble normalized variables $\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{m}}$:
$\Omega_{\mathrm{DE}}=\frac{\rho_{\mathrm{DE}}}{3H^{2}}=\frac{\rho_{\mathrm{DE}}}{\rho},\qquad\Omega_{\mathrm{m}}=\frac{\rho_{\mathrm{m}}}{3H^{2}}=\frac{\rho_{\mathrm{m}}}{\rho},$ (8)
@@
If one instead of $w_{\mathrm{DE}}(N)$ has $\Omega_{\mathrm{DE}}(N)$, then one can obtain $w_{\mathrm{DE}}(N)$ and $E_{\mathrm{DE}}^{2}(N)$ since
@@
$E_{\mathrm{DE}}^{2}$ $=\exp(-3N)\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}}}\right)=\exp(-3N)\left(\frac{1-\Omega_{\mathrm{DE}0}}{1-\Omega_{\mathrm{DE}}}\right),$ (14b)
follows from (6) and (8).
@@ 2.4 Past DE series expansions
which is indeed the case for the thawing ($w_{\infty}=-1$) and tracking ($w_{\infty}\in(-1,0)$) quintessence models. To improve the approximation we now perform series expansions in
@@
where we for brevity only consider $w_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}$ (note that $T\rightarrow 0$ when $N\rightarrow-\infty$). If one starts with a series expansion in $w_{\mathrm{DE}}$ and uses (7) and (13) one obtains an expression for $\Omega_{\mathrm{DE}}$ that subsequently can be series expanded in $T$; on the other hand, if one starts with a series expansion in $\Omega_{\mathrm{DE}}$ and inserts this into (14a) one obtains an expression for $w_{\mathrm{DE}}$ that one then can series expand in $T$. These two procedures result in the same series expansions, which can be written as [^5] The given expressions can be generalized by writing $w_{\mathrm{DE}}=w_{\infty}\left[1+\alpha T+\sigma T^{2}+\dots\right]$, which subsequently can be used to compute the series expansions for $\Omega_{\mathrm{DE}}$ and, e.g., $E_{\mathrm{DE}}^{2}$. The reason we have written $\alpha$ as $1-\gamma$ and $\sigma$ as $(\gamma-1)\beta$ is that this simplifies and is compatible with our thawing and tracking quintessence approximations, where these use the $[1/1]$ Padé expression for $\Omega_{\mathrm{DE}}$ in order to connect our approximations to the present time.
@@
$\Omega_{\mathrm{DE}}$ $=T\left(1-\gamma T+\left[1+\frac{1}{2}(\gamma-1)\left(3+\gamma+\beta\right)\right]T^{2}+\dots\right).$ (22b)
@@ 2.5 DE Padé approximants
To connect with the present time we can set $N=0$ and solve for $T=T_{0}$ in (23) in terms of the left hand side $w_{\mathrm{DE}0}$. However, since $w_{\mathrm{DE}0}$ is not an observable we will instead relate $T_{0}$ to $\Omega_{\mathrm{DE}0}$ and thereby obtain a direct connection to the observables $H_{0}$ and $\rho_{\mathrm{m}0}$ since $1-\Omega_{\mathrm{DE}0}=\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$. To obtain accurate approximations at the present time for $\Omega_{\mathrm{DE}}$ we consider the Padé approximant (use the expression (22b) up to second order in $T$)
@@
This is then inserted into (23a), which when used in (7) to obtain $f(N)$ leads to the same expression as (24) with (25) inserted. To obtain a better approximation for $\Omega_{\mathrm{DE}}(N)$ one can then go on to Padé approximants also involving $\beta$ in the expansion. [^6] These are given by $[1/2]_{\Omega_{\mathrm{DE}}}=\frac{T}{1+\gamma T-\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}$ and $[2/1]_{\Omega_{\mathrm{DE}}}=\frac{\gamma T+\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}{\gamma+\left[1+\frac{1}{2}(\gamma-1)\left(3+\beta+\gamma\right)\right]T}$. To express $T_{0}$ in terms of $\Omega_{\mathrm{DE}0}$ obtained in this manner requires solving quadratic equations in $T_{0}$, which leads to rather messy expressions. The approximation (24) turns out to be a fairly good approximation for quintessence. For this reason we will use the simple result in (25) for $T_{0}$ in (23b). This leads to that the expressions for $w_{\mathrm{DE}}$ in (23) can be written as follows: [^7] Note that one can use the series expansion in $T$ for $\Omega_{\mathrm{DE}}$ to obtain the following expression for $E_{\mathrm{DE}}^{2}$: $E_{\mathrm{DE}}^{2}\approx\Omega_{\mathrm{m}0}\exp(-3N)[1+T-(\gamma-1)T^{2}(1+{\textstyle{1\over 2}}(\gamma-1+\beta)T)+\dots]$ and use (25) for $T_{0}$. This expression can subsequently be used to obtain various Padé approximants to obtain good quintessence approximations for $E_{\mathrm{DE}}=E_{\phi}$ and then use $E_{\phi}(N)$ to compute, e.g., $w_{\mathrm{DE}}=w_{\phi}$, if one is so inclined.
@@
which result in the following present time values
@@
Note that setting $\beta=\gamma-1$ in $[1/1]_{w_{\mathrm{DE}}}(T)$ yields $[0/1]_{w_{\mathrm{DE}}}(T)$ and that setting $\gamma=1$ in both $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ results in $w_{\mathrm{DE}}=w_{\infty}$, i.e., $\gamma=1$ leads to the $w$CDM model when $w_{\infty}\in(-1,0)$ and the $\Lambda$CDM model when $w_{\infty}=-1$.
@@ 2.6 The Chevallier-Polarski-Linder parametrization
The so-called CPL parameters [25, 26] are often used in cosmology [^9] For a recent discussion of the CPL parametrization, see Wolf & Ferreira [33]. and parametrize $w_{\mathrm{DE}}$ according to
@@
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{\Omega_{\mathrm{m}0}^{2}},$ (33a)
@@
One can regard (31) as a series expansion in $e^{N}$, but this is inconsistent with the past matter dominant requirement $\Omega_{\mathrm{DE}}\rightarrow 0$ and that one of the CPL parameters can be interpreted as $w_{\infty}$, since this naturally leads to an expansion in $e^{-3w_{\infty}N}$, which is particularly pertinent as observations probe increasingly large redshift. CPL thus mixes this asymptotically motivated function (arising from the $w_{\infty}$ part in (31)) in $\Omega_{\mathrm{CPL}}$ and $E_{\mathrm{CPL}}^{2}$ with another function that comes from the $w_{a}$ part in (31). Due to this mixture of functions, especially the exponential part involving $\frac{z}{1+z}$, the CPL parametrization arguably does not lead to simpler expressions for observables than our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based results, and it certainly is not a good approximation for thawing and tracking quintessence.
@@ 3.1 Thawing quintessence
To obtain a dynamical system that is useful for deriving expressions for $\gamma$ and $\beta$ for thawing quintessence [^10] The concepts of thawing and freezing were defined by Caldwell and Linder (2005) [34] as follows: thawing is characterized by $w_{\varphi}\approx-1$ where $w_{\varphi}$ subsequently grows, i.e. $w_{\varphi}^{\prime}>0$, while $w_{\varphi}>-1$ and $w_{\varphi}^{\prime}<0$ holds for freezing. we introduce the following dimensionless bounded variable [^11] The variable $\Sigma_{\phi}$ was first introduced by Coley et al. (1997) [35] and Copeland et al. (1998) [36] whose $x$ is $\Sigma_{\varphi}$. Since then, $\Sigma_{\phi}$ (or $\phi^{\prime}$) is commonly used to describe scalar fields in cosmology, see, e.g., Urena-Lopez (2012) [37], equation (2.3), Tsujikawa (2013) [28], equation (16) and Alho and Uggla (2015) [38], equation (8). The reason for using the notation $\Sigma$ for the kernel is because $\Sigma_{\phi}$ plays a similar role as Hubble-normalized shear, which is typically denoted with the kernel $\Sigma$, see e.g. [39].
@@
as follows from the definitions. In addition, the (non-linear) Klein-Gordon equation $\ddot{\varphi}=-3H\dot{\varphi}-V_{,\varphi}$, expressed in $e$-fold time and $\Sigma_{\phi}$, is needed. This leads to the dynamical system